OpenStudy (anonymous):

Which of the following ordered pairs is the solution to the system of linear equations x+y=14 2x+y= 24 A.(7, 7) B.(3, 11) C.(6, 12) D.(10, 4)

Parth (parthkohli):

Subtract first from second equation.

OpenStudy (anonymous):

then what do i do

OpenStudy (shane_b):

Show the subtraction...if you do the work, you'll see what x is and there's only one answer with that x value.

OpenStudy (anonymous):

im not sure wAT TO SUBTRACT

OpenStudy (shane_b):

As @ParthKohli said, do this: 2x+y= 24 - (x+y)=14 ------------- =?

OpenStudy (anonymous):

10

OpenStudy (shane_b):

Yes, x = 10...so y must be 4 (look at the first equation to see that).

OpenStudy (anonymous):

10,4 thanks alot